The airplane flies 1080 miles against the wind in 4 hours and the return trip takes 3 hours. Find the speed of the airplane with no wind and the speed of the wind.
I know that it is d=r*t
flying against the wind
(A-W) *4= 1080
A-W= 1080/4= 270
flying with the wind
(A+W)* 3= 1080
A+W=1080/4= 360
Is this correct? What do I do next?

Answer by Alan3354(69443)

The airplane flies 1080 miles against the wind in 4 hours and the return trip takes 3 hours. Find the speed of the airplane with no wind and the speed of the wind.
I know that it is d=r*t
flying against the wind
(A-W) *4= 1080
A-W= 1080/4= 270
flying with the wind
(A+W)* 3= 1080
A+W=1080/4= 360 ******** 1080/3 = 360
Is this correct? What do I do next?
--------------------
It's correct except for the noted typo
---
a - w = 270
a + w = 360
-------------- Add
2a = 630
a = 315
w = 45